The Millers plan to serve a two-pound banana split for a dinner party dessert. The recipe calls for two and a half times as much ice cream as syrup and one and two thirds times as much banana as syrup plus three and seven twelfths ounces of whipped cream and cherries. How many ounces of ice cream should be used?
Answer by richwmiller(17219)

2.5s=c,
b=(5/3)*(s+43/12),
w+h=43/12,
c+s+w+h=32
bananas b=19.504, ice cream c=20.2976,
syrup s=8.11905, whipped cream and cherries w+h=3.58333